OpenStudy (jake124):

5/10 - 3/10 A. 1/5 B. 8/10 C. 42/5 D. 1/10

OpenStudy (lowkey.s):

any ideas?

OpenStudy (misssmartiez):

5 - 3 = 2.

OpenStudy (lowkey.s):

NO

OpenStudy (misssmartiez):

2/10 WITH THE GREATEST COMMON WHAT?

OpenStudy (lowkey.s):

thats not how u do it

OpenStudy (lowkey.s):

DONT PUT CAP ON ME

OpenStudy (misssmartiez):

Yes indeed it is. Since our denominators are the same, we do not need to attend those values.

OpenStudy (misssmartiez):

First of all, now we subtract the numerators, keeping the denominators the same.

OpenStudy (lowkey.s):

its LCD

OpenStudy (misssmartiez):

Which is 2.

OpenStudy (jake124):

i think its b

OpenStudy (lowkey.s):

it A

OpenStudy (misssmartiez):

No dear.

OpenStudy (lowkey.s):

yes dear

OpenStudy (misssmartiez):

Stop Lowkey, do not give direct answers.

OpenStudy (lowkey.s):

It's LCD

OpenStudy (lowkey.s):

well im not

OpenStudy (misssmartiez):

Anyway, as I said, we need to figure out what number the two factors have in common. The greatest for the best simplification.

OpenStudy (lowkey.s):

No its not its least common dominator

OpenStudy (jake124):

thanks can u help me with more ?

OpenStudy (misssmartiez):

The highest it could go without surpassing 2, is 2 itself. 2 / 10 = 5 2 / 2 = 1 Therefore its A, 1/5

OpenStudy (lowkey.s):

so you do LCD

OpenStudy (lowkey.s):

not GCD

OpenStudy (misssmartiez):

@Lowkey.S I suggest you stop yelling, interpreting, and acting as such. I am merely explaining the conclusion.

OpenStudy (misssmartiez):

It would be GCD because 2's greatest factor that corresponds and can go INTO 10, is 2.

OpenStudy (lowkey.s):

you always still my question! and you probably know who i am

OpenStudy (jake124):

what calulator should i use to calculate this stuff

OpenStudy (misssmartiez):

I steal your question? Dear, I always medal you Lowkey.

OpenStudy (misssmartiez):

Or simply give my output, if it felt as if, I apologize, but I would ever so still acknowledge you and medal you first.

OpenStudy (misssmartiez):

For now jake, you don't necessarily need a calculator to do this process. It is simple if you think about it :).

OpenStudy (lowkey.s):

okay then